given functions f and g, determine the domain of f+g
f(x)=3x^2 -9 g(x)=2x^3 -10

Answer by nyc_function(2741)

f + g = f(x) + g(x).
f + g = 3x^2 - 9 + 2x^3 - 10
f + g = 2x^3 + 3x^2 - 19
The domain of f + g is ALL REAL NUMBERS because x can be safely replaced by any real number in the function.